The total human population of the planet Earth is something in the vicinity of 7 billion. If you choose one human being at random from across the entire planet, each individual on the planet is equally likely to be chosen. Corresponding probabilites are impractically small. Use this information to answer the following questions This human population of the entire planet Earth can be divided roughly into thirds: About 1/3 of the population is adult (ages 18 -65) men; about 1/3 is adult women; and about 1/3 is non-adults (children and older adults of either gender). Fractions (or proportions or percents) of the population can also be probabilities. If you choose one human being at random from the entire planet Earth, individual probabilities are (still) vanishingly small, but the fraction in any given group can be used as the probability of choosing someone from that group.

If you pick one human being from the population of the planet Earth, what is the probability that you get an adult of either gender?

FRACTION answer.
Answer #1

Pr(Adult of either gender) = P ( adult men ) + P( adult women ) = 1/3 + 1/3 = 2/3

So required probability is 2/3 .

Vote if this is useful.